### v1

@@ 1 Introduction
To avoid degeneracy between the natural variation of light-curve width and time dilation, Foley et al. (2005) and Blondin et al. (2008) observed time dilation in the evolution of spectral features of high-$z$ Type Ia supernovae (SNe Ia). The former found inconsistency with no time dilation at the 99% confidence level and the latter finding $b=0.97\pm 0.10$. Most recently, Lewis & Brewer (2023) inferred $b=1.28^{+0.28}_{-0.29}$ using the variability of 190 quasars out to $z\sim 4$. Despite these successes, there remains continued discussion of hybrid or static-universe models such as Tired Light (Zwicky, 1929; Gupta, 2023) that do not predict expansion-induced time dilation.
@@ 3 Data
We exclusively use the data of the 1635 type Ia supernovae measured by the Dark Energy Survey Supernova Program (DES Collaboration et al., 2024). The DECam instrument on the 4m Blanco telescope at the Cerro Tololo Inter-American Observatory (Flaugher et al., 2015) observed most of the photometrically classified SN Ia candidates in the $g$, $r$, $i$, and $z$ bands according to the criteria in Smith et al. (2020). The flux is determined by difference imaging (Kessler et al., 2015). High-redshift SNe Ia typically show negligible flux in the ultraviolet wavelength region, so $g$ and $r$ band light curves are only useful for SNe at $z\lesssim 0.4$ and $z\lesssim 0.85$ respectively (see DES Collaboration et al., 2024, Fig. 2). The SALT3 (Kenworthy et al., 2021) template fits with a cadence of 2-day time sampling in each band were available for each SN candidate. We use these fits to estimate the peak flux of each supernova so we can normalize the observed flux values; the time relative to the $B$-band maximum in the SALT3 fit for each light curve is also used to define the time since peak brightness measurement in our analysis. We otherwise discard the SALT supernova information.
@@ 4.1 Reference curve construction
After populating the reference curve with data points we then normalise the photometry in flux; as the curve is populated with the data of several SNe at different redshifts, the curve must be homogeneous in flux. To do this, we utilised the peak flux in the SALT3 model light curves provided for each SNe. The data in each constituent curve is normalised by this value before being added to the reference. For convenience we also use the time of peak brightness given by SALT3 as the reference point about which to stretch the light curves (see equation 5). These are the only uses of SALT3 information and we expect that the same time dilation signal would be obtained in the data with any other consistent normalisation measures.
@@ 4.2 First measure of time dilation: minimising scatter in the reference curve
If there was no time dilation we would expect the minimum dispersion in the reference curve to be at $b\sim 0$ (i.e. no time scaling) in Fig. 5. The fact that we find $b\approx 1$ is evidence for time dilation of the expected form. This a rough but completely model independent measure of time dilation and it is the paper’s first main result.
@@ 4.3 Second measure of time dilation: Finding each light curve width
Figure 6: We show the reference curve construction and subsequent target SN fit for 3 SNe at redshifts $z\simeq 0.22$, $z\simeq 0.43$, and $z\simeq 1.02$ and in fitting bands $r$, $i$, and $z$ respectively (in descending order). The left plots show the allowed ranges for reference curve SN sampling given the target redshift (and $\delta=2^{-4}$). The vertical line of dots is plotted at the target SN redshift, with each dot representing the redshift of a DES supernova (vertical axis). The dots that fall in the narrow coloured bands are the SNe that make up the reference population, as those data all share approximately the same rest-frame wavelength in their respective bands. The right plots show the constructed $(1+z)$ time-scaled reference curve (small coloured points) with respect to the target SN photometry (blue points) and subsequent target photometry scaled on the time axis to fit the reference (best-fit widths of 1.42, 1.49, and 2.17 respectively). Due to the statistics associated with such large reference curve populations, the contribution of any individual reference point uncertainty to the overall reference curve uncertainty is negligible and not plotted; the uncertainty in the target data has a much higher contribution to the uncertainty in the fitting.
@@
We note that this method of fitting is not fundamentally limited to target SN data with pre-peak brightness observations in each band. Given enough target data (on the order of several well space points in time), the mapping of this data to the corresponding reference curve phases is unique regardless of whether pre-peak data is available.
@@ Appendix A Stretch drift with redshift
where $\mathcal{N}(\mu,\sigma^{2})$ is a normal distribution with mean $\mu$ and variance $\sigma^{2}$ and the values of the parameters were $(a,\mu_{1},\mu_{2},\sigma_{1},\sigma_{2},K)=(0.51,0.37,-1.22,0.61,0.56,0.87)$; and the fraction of young supernovae in the population is given by,
@@
The distribution given by equation (8) is shown in the upper panel of Fig. 10 for several redshifts, where the vertical dashed lines show the resulting change in the mean $x_{1}$. The relationship between $x_{1}$ and the stretch of the supernova is given by (Guy et al., 2007),
@@
Nevertheless, in the DES-SN5YR data there is no indication that the mean stretch parameter $x_{1}$ changes with redshift, see Fig. 11.
@@
Figure 11: The distribution of stretch in the DES-SN5YR data as a function of redshift (calculated from the SALT3 fitted $x_{1}$ values using equation 9), with $x_{1}$ shown on the right axis. Fitting a straight line to this distribution shows no significant trend in the stretch with redshift.
@@ Appendix C Null test — no de-redshifting of reference light curves
1 School of Mathematics and Physics, The University of Queensland, QLD 4072, Australia

### v2

@@ 1 Introduction
To avoid degeneracy between the natural variation of light-curve width and time dilation, the evolution of spectral features of high-$z$ SNe Ia was used to measure time dilation by Riess et al. (1997), Foley et al. (2005), and Blondin et al. (2008). The first two of these studies found inconsistency with no time dilation at the 96.4% confidence level (for SN1996bj) and 99% (for SN1997ex) respectively, and the third finding $b=0.97\pm 0.10$. Most recently, Lewis & Brewer (2023) inferred $b=1.28^{+0.28}_{-0.29}$ using the variability of 190 quasars out to $z\sim 4$. Despite these successes, there remains continued discussion of hybrid or static-universe models such as Tired Light (Zwicky, 1929; Gupta, 2023) that do not predict expansion-induced time dilation.
@@ 3 Data
We exclusively use the data of the 1635 type Ia supernovae measured by the Dark Energy Survey Supernova Program (DES Collaboration et al., 2024). The DECam instrument on the 4m Blanco telescope at the Cerro Tololo Inter-American Observatory (Flaugher et al., 2015) observed most of the photometrically classified SN Ia candidates in the $g$, $r$, $i$, and $z$ bands according to the criteria in Smith et al. (2020). The flux is determined by difference imaging (Kessler et al., 2015). High-redshift SNe Ia typically show negligible flux in the ultraviolet wavelength region, so $g$ and $r$ band light curves are only useful for SNe at $z\lesssim 0.4$ and $z\lesssim 0.85$ respectively (see DES Collaboration et al., 2024, Fig. 2). The SALT3 (Kenworthy et al., 2021) template fits with a cadence of 2-day time sampling in each band were available for each SN candidate. We use these fits only to estimate the peak flux of each supernova and the associated time of peak flux (given by the $B$-band maximum time in the observer frame) so that we can normalize the light curves in brightness and in time relative to the peak brightness. We otherwise discard the SALT supernova information.
@@ 4.1 Reference curve construction
After populating the reference curve with data points we then normalise the photometry in flux; as the curve is populated with the data of several SNe at different redshifts, the curve must be homogeneous in flux. To do this, we utilised the peak flux in the SALT3 model light curves provided for each SNe, which is independent of the time dilation signal. The data in each constituent curve is normalised by this value before being added to the reference. For convenience we also use the time of peak brightness given by SALT3 as the reference point about which to stretch the light curves (see equation 5), and this parameter is also in the observer frame of reference. These are the only uses of SALT3 information and we expect that the same time dilation signal would be obtained in the data with any other consistent normalisation measures.
@@ 4.2 First measure of time dilation: minimising scatter in the reference curve
If there was no time dilation we would expect the minimum dispersion in the reference curve to be at $b\sim 0$ (i.e. no time scaling) in Fig. 5. The fact that we find $b\approx 1$ is evidence for time dilation of the expected form. This is a rough but completely model independent measure of time dilation and it is the paper’s first main result.
@@ 4.3 Second measure of time dilation: Finding each light curve width
Figure 6: We show the reference curve construction and subsequent target SN fit for 3 SNe at redshifts $z\simeq 0.22$, $z\simeq 0.43$, and $z\simeq 1.02$ and in fitting bands $r$, $i$, and $z$ respectively (in descending order). The left plots show the allowed ranges for reference curve SN sampling given the target redshift (and $\delta=2^{-4}$). The vertical line of dots is plotted at the target SN redshift, with each dot representing the redshift of a DES supernova (vertical axis). The dots that fall in the narrow coloured bands are the SNe that make up the reference population, as those data all share approximately the same rest-frame wavelength in their respective bands; boxes are plotted surrounding these regions to more clearly indicate the light curve colour samples that make up the reference curve. The right plots show the constructed $(1+z)$ time-scaled reference curve (small coloured points) with respect to the target SN photometry (blue points) and subsequent target photometry scaled on the time axis to fit the reference (best-fit widths of 1.42, 1.49, and 2.17 respectively). Due to the statistics associated with such large reference curve populations, the contribution of any individual reference point uncertainty to the overall reference curve uncertainty is negligible and not plotted; the uncertainty in the target data has a much higher contribution to the uncertainty in the fitting.
@@
We note that this method of fitting is not fundamentally limited to target SN data with pre-peak brightness observations in each band. Given enough target data (on the order of several well-spaced points in time), the mapping of this data to the corresponding reference curve phases is unique regardless of whether pre-peak data is available.
@@ 6 Conclusions
Contribution Statement
RMTW performed the main analysis and drafted the manuscript; TMD supervised the project, performed the redshift drift analysis, and wrote/edited sections of the paper; GFL assisted in testing the code and developing ideas; PS, CL were internal reviewers and RK the final reader. JDV, HTD, KG, and AM gave feedback on the draft manuscript; AM, DB, RK advised on the data and analysis; the remaining authors have made contributions to this paper that include, but are not limited to, the construction of DECam and other aspects of collecting the data; data processing and calibration; developing broadly used methods, codes, and simulations; running the pipelines and validation tests; and promoting the science analysis.
@@ Appendix A Stretch drift with redshift
where $\mathcal{N}(\mu,\sigma^{2})$ is a normal distribution with mean $\mu$ and variance $\sigma^{2}$ and the values of the parameters were $(a,\mu_{1},\mu_{2},\sigma_{1},\sigma_{2},K)=(0.51,0.37,-1.22,0.61,0.56,0.87)$. The fraction of young supernovae in the population is given by,
@@
where the parameter $K$ is related to the star formation rate within galaxies (Rigault et al., 2013). The distribution given by equation (8) is shown in the upper panel of Fig. 10 for several redshifts, where the vertical dashed lines show the resulting change in the mean $x_{1}$. The relationship between $x_{1}$ and the stretch of the supernova is given by (Guy et al., 2007),
@@
Nevertheless, in the DES-SN5YR data there is no indication that the mean stretch parameter $x_{1}$ changes with redshift, see Fig. 11. The stretch values shown in this plot are from the SALT3 fits to the light curves and are in the rest frame of the supernova (i.e. corrected for time dilation).
@@
Figure 11: The distribution of stretch in the DES-SN5YR data as a function of redshift (calculated from the SALT3 fitted $x_{1}$ values in the SN rest frame using equation 9), with $x_{1}$ shown on the right axis. Fitting a straight line to this distribution shows no significant trend in the stretch with redshift.
@@ Appendix C Null test — no de-redshifting of reference light curves
1 School of Mathematics and Physics, University of Queensland, Brisbane, QLD 4072, Australia
@@ 13 Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK
17 Instituto de Física Teórica, Universidade Estadual Paulista, Sáo Paulo, Brazil
@@ 52 Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA
53 ASTRAVEO LLC, PO Box 1668, MA 01931, USA
54 Applied Materials Inc., 35 Dory Road, Gloucester, MA 01930, USA